OpenStudy (anonymous):

Thanks for helping! I was given a pre calc problem which I simplified to: (tanx-2)(tanx-1)=0 The question asks me to solve from there. However, it doesn't ask me to solve in terms of 0 to 2pi, it asks for all solutions. I know that I have to add an extension once I find the final solution, but am having a hard time figuring out how to do so. I know it looks something like+n*pi*k or something like that. Thank You!

OpenStudy (anonymous):

@Kainui @jim_thompson5910 @amistre64 @ganeshie8 @zepdrix @campbell_st @Loser66

OpenStudy (anonymous):

@Luigi0210 @robtobey @sammixboo @Zarkon @pooja195 @radar

zepdrix (zepdrix):

Hmm the 2 is going to cause a bit of a problem :) Not a nice friendly angle there. But the 1 is ok.

OpenStudy (anonymous):

I think the one is pi/4?

zepdrix (zepdrix):

Applying our Zero-Factor Property,\[\Large\rm \tan x-1=0\]Add 1 to each side,\[\Large\rm \tan x=1\] Ok good, yes this corresponds to pi/4 in the first quadrant.

OpenStudy (anonymous):

since it is for all solutions, should I add the extension? Or is there another thing I need to do?

zepdrix (zepdrix):

\[\Large\rm x=\frac{\pi}{4}+k \pi, \quad k=0,\pm1,\pm2,...\]Tangent is a little different than sine and cosine. It is period in `pi` not in `2pi`. So we want to allow multiples of pi to be added on and give us the same value through tangent.

zepdrix (zepdrix):

periodic* blah

OpenStudy (anonymous):

So I'd just say pi/4+kpi and add nEz to the right?

zepdrix (zepdrix):

\(\Large\rm k\in\mathbb{Z}\) ya that works out nicely I guess :)

OpenStudy (anonymous):

cool and I don't need to add 5pi/4+kpi? As it is another solution

zepdrix (zepdrix):

5pi/4 = pi/4 + 1*pi so it's already included in our set of solutions with the k

zepdrix (zepdrix):

we just need to deal with this ugly tanx-2 now :d

OpenStudy (anonymous):

lol, I simplified it to 7pi/20

OpenStudy (anonymous):

I think that sounds right, but am not sure

zepdrix (zepdrix):

Woah I'm not really sure 0_O I was just going to be lazy and call the angle \(\Large\rm \arctan(2)\)

OpenStudy (anonymous):

I asked my teacher, but he said I needed to go further

zepdrix (zepdrix):

mmm ok thinking :)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

I found the radian in decimals and divided by pi and converted that into a fraction

OpenStudy (anonymous):

If that is the correct process

OpenStudy (anonymous):

and multiplied pi of course :)

zepdrix (zepdrix):

arctan(2) = 1.1071487... divide by pi gives: \(\Large\rm 0.352416382\pi\) then what did you do? 0_o how did you convert to fraction?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

about 7/20

OpenStudy (anonymous):

is that right? I'm not too good at this ha ha

OpenStudy (anonymous):

I converted with some calculator

OpenStudy (anonymous):

which gave me 7/20, which is about right

OpenStudy (anonymous):

I rounded it for the calc

zepdrix (zepdrix):

im not sure how to get an exact answer :d its look like it's some irrational multiple of pi, like some weird square root maybe. so ya approximating is probably the right way to go. I would leave it as a decimal though maybe. \[\Large\rm x=0.352\pi+k \pi\]\[\Large\rm x=(0.352+k)\pi,\qquad k\in\mathbb{Z}\]

zepdrix (zepdrix):

Maybe you guys have some fun technique for finding arctan(2), maybe my brain is a lil rusty hehe

OpenStudy (anonymous):

thanks for the help! I'll give you a medal

zepdrix (zepdrix):

np c:

OpenStudy (anonymous):

:)